buttercup888:

Part A: The sun produces 3.9 • 1033 ergs of radiant energy per second. How many ergs of radiant energy does the sun produce in 3.25 • 103 seconds? (5 points) Part B: Which is the more reasonable measurement of the distance between the tracks on a railroad: 1.435 • 10–3 mm or 1.435 • 103 mm? Justify your answer.

Vocaloid:

part A: just multiply ergs per second * number of seconds part B: 1.435 * 10^-3 is VERY small so it's not logical; choose the other one

buttercup888:

I dont understand

Vocaloid:

for part A: you are given ergs per second (3.9 • 10^33) and the number of seconds (3.25 • 10^3) just multiply these to get the answer

buttercup888:

so 3.9*10^3=3900 and 3.25*10^3=3250 so then 3900*3250. it that right

Vocaloid:

I think it says 3.9*10^33 not 3.9*10^3 ^^"

Vocaloid:

(3.9 • 10^33)(3.25 • 10^3) = ?

Vocaloid:

hint: to multiply 10^33 * 10^3 just add the exponents and raise 10 to that power

buttercup888:

Okay.

buttercup888:

so then 3.9*10^33=39000 and 3.25*10^33=32500

Vocaloid:

nope

Vocaloid:

again, take 10^33 and 10^3 add the exponents together and let me know what you get

buttercup888:

cant you just give me the answer pleeeeeeaaaasssseee

Vocaloid:

the exponents are 33 and 3 add these together.

buttercup888:

36

buttercup888:

so 10^36

Vocaloid:

good now we just multiply the other parts 3.9*3.25 = ?

buttercup888:

12.675

Vocaloid:

good now we can write our answer as 12.675 * 10^36 we can re-write this in scientific notation as 1.2675 * 10^37 ergs = your answer

buttercup888:

Thank you so much

Vocaloid:

and for part B, to reiterate, it would be 1.435 • 10^3 mm b/c the other one is too small to be reasonable

buttercup888:

Okay thanks